OpenStudy (anonymous):

Find the equation of the line whose slope is 2 and which passes through the point (-2, 4).

OpenStudy (anonymous):

with the given information you can plug all that into \[y - y_{1}=m(x-x_{1})\] and solve for y and then you will have a pretty \[y = mx + b\] equation

OpenStudy (anonymous):

so i will start you out \[y-4 = 2(y+2)\] Do you know what to do from here?